OpenStudy (anonymous):

Help! -7a^2 (3a^3 - 4a^5)

OpenStudy (cggurumanjunath):

-21a^5+28a^7

OpenStudy (cggurumanjunath):

-7a^2*3*a^3 =-7*3*a^(2+3)

OpenStudy (cggurumanjunath):

-7a^2(-4a^5)=-7*4*a^(2+5)

OpenStudy (cggurumanjunath):

add the two ! we get the first (final) answer